OpenStudy (tthreet):

It is weigh-in time for the local under 85 kg rugby team. The bathroom scale used to assess eligibility can be described by Hooke\'s law, which is depressed 0.67 cm for its maximum load of 115 kg. What is the spring\'s effective spring constant?

OpenStudy (zpupster):

F=-kx where m=115kg, x=.65*10^-2m k=-F/x=mg/x=??